Circular No. 6807

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVAE L. Germany reports, on behalf of the Mount Stromlo Abell Cluster

Supernova Search Team (cf. IAUC 6639), their discovery of four apparent supernovae on V and R CCD images (limiting mag V = 21.5) taken by S. Chan at the Mount Stromlo 1.27-m telescope (+ Macho Camera):

SN  1997 UT  α_{2000}  δ_{2000}  V  Offset
1997fb Dec. 31 5^{h}01^{m}14.^{s}60 −38^{o}38′12′′.7 18.6 1′′.3 W, 0′′.4 N
1997fc Dec. 31 4 59 58.81 −38 55 04.3 20.5 4′′.9 E, 1′′.9 N
1997fd Dec. 28 3 41 15.65 −53 29 30.8 21.2 7′′.1 E, 1′′.3 S
1997fe Dec. 31 4 33 07.64 −61 26 19.3 19.5 5′′.7 E, 6′′.2 S

Confirmation images (limiting mag R = 21) for all four objects were obtained by S. Sabine at the same telescope on 1998 Jan. 12. Previous images not showing each supernova were taken on the following dates (with galaxy cluster membership noted parenthetically): 1997fb and 1997fc, Dec. 19 (Abell 3301); 1997fd, Dec. 12 (Abell 3158); 1997fe, Nov. 7 (Abell 3266).

PKS 2255–282 D. Bertsch, Goddard Space Flight Center, NASA, on behalf of the

EGRET Collaboration, writes: “The EGRET instrument on the Compton Gamma Ray Observatory (CGRO) has detected a γ-ray source at energies above 100 MeV (significance 6σ) located at α = 22^{h}55^{m}, δ = −28^{o}.2 (equinox 2000.0). The source’s 50-percent uncertainty contour includes PKS 2255– 282, a flat-spectrum radio quasar of visual mag 16.7 and redshift 0.9, and which has catalogued radio fluxes of 1.8 Jy at 5 GHz and 3 Jy at 31 GHz. A preliminary analysis indicates that the photon intensity is 3–4 × 10^{−6} photons cm^{−2} s^{−1}, integrated above 100 MeV (from data obtained between 1997 Dec. 30 and 1998 Jan. 3). EGRET had not detected this source previously, so that a strong flare is indicated. The data will be released to the public archive at the CGRO Science Support Center, as soon as initial processing is completed. A preliminary view of the likelihood map is available

at the Web site http://cossc.gsfc.nasa.gov/cossc/2255 282.gif.”

GRB 971227 Further to their item on IAUC 6803, S. A. Ilovaisky reports that the

CCD image of the GRB 971227 field noted therein can be found at URL

http://www.obs-hp.fr/~ilovaisky/grb.html.

1998 January 15  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green